Home‎ > ‎Reports‎ > ‎Crew 109‎ > ‎

day02


Crew #109 Log Book for December 18, 2011

Photos of the day





Commander's Check-In Report

Date and time: 18 December 2011

Written by Nicky De Munster, Crew 109, Commander:
 
Crew Physical Status: some headache, good
 
Time Departed/Returned from EVA:

EVA 1: departed: 8:28 am / Returned 10:05 am
EVA 2: departed: 2:15 pm / Returned 4:10 pm
 
Brief Narrative of Field Mission Results:
 
EVA Data/Interpretations:
 
Engineering/Hab Maintenance:
 
Report Transmission Schedule:
 
Plans for Tomorrow: 2 EVA's, analyse soil samples, water samples
 
Inventory: Food inventory OK
 
Support Requested: none

Water level is low, must be supplied.
Gasoline level is low, just enough for another two days. 

Miscellaneous: N/A


Commander's Report

Date and time: 18 december 2011

Written by Nicky De Munster, Crew 109, Commander:

Text:

Yesterday after the crew had left, we had a first meeting on the schedule for today. Because the weather was very good and the predictions were good too, we decided to startimmediately with the EVA's.

Karon made ​​a plan with kitchen and EVA rotations. While a crew is on EVA is, the other works in the hab and will do the cooking.

We started the food study today by weighing the crew. The first EVA departed already at 8:28 am. The fog on the helmet was a major problem. Radio contact was very poor. Karon took some soil samples to analyse. David took some rock samples.

The second EVA had no fogging problems with the helmets because of the higher temperatures. Radio contact was again very bad.
I took some water samples to analyse in the lab tomorrow and also some soil samples to start the seeds project. Victoria took soil samples to analyse in the lab.

Aster tested successfully a new helmet camera.

The mood is good.


Science Report

Date and time: 18 December 2011, 18:00 

Written by David Kutai Weiss, Crew Geologist & Executive Officer, 
Karon Wynne, Health & and Safety Office and Crew Biologist, Victoria Grudzinski, Crew Biologist


Biology Report:

Victoria and Karon began the first day of the mission determining the logistics of how to collect biological samples using aseptic techniques during EVAs. We took some samples in sites where the soil was moist, as we plan to plate these samples to look for extremophiles. The objective of this study is to determine if there are any bacteria capable of living in the harsh, cold conditions surrounding MDRS. Karon collected 2 samples.

The biology team also took advantage of the recent snowfall and collected samples to test the pH, oxygen level and conductivity of the water. We want to determine the logistics of running such tests as the properties of water found on a true Manned mission to Mars would have very important implications.


Geology Report

Project Relevance: Any comprehensive human-engaged Mars mission proposal, suggests the use of robotic exploration of the surface of Mars to augment human activity where humans have neither the capacity nor the time to explore. Robotic exploration combined with human field work during a manned mission to Mars is important in resource acquisition; namely in locating water. Because a rover during such a mission would be controlled by a ground crew on the surface of Mars and would not have a communications time delay, it would be able to cover much more ground than any previous Mars rovers. Because more ground will be covered, it is important to know where the rover could physically travel, which is constrained by surface geology and terrain gradient.

While the water for an initial human occupation of Mars could be brought in the form of hydrogen, and acquiring the oxygen from carbon dioxide in the atmosphere, long term water supply is a concern for future missions seeking a prolonged stay on Mars. One proposed idea for generating water on Mars is by acquiring water from Mars’ regolith through heating.

Because the device which collects water from the Martian soils may be a rover-like device with limitations on where it can travel based on local terrain, it is important to link the available routes with high-water-content soils. On Mars, generating a map of the area surrounding the habitat that shows which high water content soils are easiest to reach by rover will be imperative in water collection. Mapping out viable rover routes and soil properties will allow us to carry out robotic and human exploration for water in the soils of Utah, and thus the soils of Mars.


Proposed Project: 

During a Mars Desert Research Station (MDRS) rotation I will combine satellite imagery and ground based observations to create a map of potential rover exploration routes for the area surrounding the MDRS habitat, including routes to areas of high water content. My project seeks to simulate data sets available for the Martian surface in order to perform a comprehensive study of the feasibility for this to be carried out on the surface of Mars. Cone penetrometer data -a blow-count test, site photography, and sample collection for sieving and water content measurement can be conducted on Extra Vehicular Activities (EVAs) to produce detailed maps with valuable information and the most effective rover routes.

Today is day 1 here on “Mars”, and David spent the previous night generating a map of the surrounding area from satellite data. He added a layer to my map which plots soil types, and intends to travel to each soil-area to measure their soil strength and water content. This afternoon, David identified 12 prime sites* to do my soil measurements and plotted them on a map for future EVA’s. Tomorrow, he intends to begin measuring the soil strength with a cone penetrometer blowcount test, and measuring water content by taking core samples. Later, he will place the core samples in an oven to degass the water from them, from which the water content of the soil can be inferred.

EVA 1&2 Report

Date and Time: 18 Decembet 2011 18:30 

Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

EVA 1:
Time: 8:30 am  - 10:15 am.
Crew Members: David, Leandro, and Karon
Site Location: North and west of Hab
Transit Mode: ATV, on foot.
Objectives: General exploration

Results:

Route spanned 7 miles, we accustomed ourselves to the environment.
Karon took soil sample at stop 1.
Frozen pond observed, marked on GPS by Leandro.

EVA 2:
Time: 2:15 pm - 4:00 pm.
Crew Members: Nicky, Aster, and Victoria
Site Location: North of Hab
Transit Mode: ATV, on foot.
Objectives: General exploration, initial sample acquisition.

Results:

Route spanned 9km.
Nicky took soil, water, and snow samples
Victoria took snow samples at stop 2 and final stop
Aster tested a helmet camera to document future EVA’s.

EVA 1 & 2 Lessons learned: 

Helmets fog up less later in day during higher temperatures, but in morning the helmets fog up regardless of soaping the visors.  Radio communication during EVA is virtually impossible because when headsets are on, the radio still only receives auditory input from main radio mic, not headset mic – the helmets reduce noise so much that it is impossible to hear anyone.  Radio communication between Hab and EVA is also impossible at any distance not proximal to hab. 

Hand signals are the only viable form of communication during EVA.

Engineering Report

Date and time: 18 December 2011, 18:17

Written by Leandro G. Barajas, Crew 109, Chief Engineer:

Generator/Electricity:
Did you have a power outage: No

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): 0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): Empty

Notes/Comments: All 3 ATVs fuel tanks are full for tomorrow's EVA however there are only 4 galons of gas left.


Oil:
Kitty Oil Quantity (# of Quarts in storage): N/A
Honey Oil Quantity (# of Quarts in storage): N/A
ATV Oil Quantity (# of Quarts in storage): 1.3
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 14
Trailer Potable Water Tank Level (inches from bottom): 12
Hab Potable Water Tank Level (inches from bottom): 8
Potable Water Meter Reading: 477840
In to GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A
Greenhab Notes/Comments: N/A

Observatory:
Telescope Used? (Yes/No): NO
Observatory Notes/Comments: N/A
Transportation:


Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: N/A
Notes/Comments on Hab Car: N/A

ATV 1 (Opportunity):
Used (yes/no): YES
Oil Checked (yes/no): NO
Fuel Consumed: 1 Gal
Tire Status: Good
Comments:

ATV 2 (Spirit):
Used (yes/no): YES
Oil Checked (yes/no): NO
Fuel Consumed: 1 Gal
Tire Status: Good
Comments:

ATV 3 (Viking I):
Used (yes/no): YES
Oil Checked (yes/no): NO
Fuel Consumed: 1 Gal
Tire Status: Good
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 70

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 58

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:
Everything OK except that observatory cameras are offline. Observatory Breaker is Off. Should this be kept off?

General Engineering Notes/Comments:
Heat Tapes Operating 24/7 (Both Lights On)
Mars Surface Suits OK except for number 4 which was being rebuild by Crew 108. The main power switch is missing.
Requesting permission to continue the rebuild process and install a spare switch.
Only 3 Radios operational at this time. However they can not be effectively used with the suits. Most headsets are not working.
Most Batteries and chargers are not operational.

Journalist Report

Date and time: 18 December 2011 18:30

Written by : Aster Stein

Text:

The whole crew not only survived the first night at the Hab, but everyone was quite positive about it. There were no sleeping problems reported. Nicky and Aster were on cooking duty today, and had started the breadmaker before going to bed. The bread wasn't removed from the machine for a few hours after it was done. The bread turned out OK, but strangely "compact". Maybe this is due to the amount of yeast used, or the extended time in the machine.

Our breakfast consisted of bread with peanut butter/jelly, a small chunk of cheese and some re-hydrated milk or lemonade. We have established a schedule for 2 EVA's per day. Today, the AM-EVA was conducted by Karon, Leandro and David. The suiting up was very exiting for everyone, and soon after our first EVA-crew disapeared behind the martian hills. Radio contact was really difficult, among the EVA-crewmembers and impossible between EVA and the Hab. They all returned safely, but the helmets were completely fogged up. In the meantime, Nicky and Aster started making an inventory of the chaos of the kitchen cabinets.

Nicky brewed up a lovely "add-water" gumbo soup for lunch. Aster prepared a big bowl of "add-water" chili for diner. Stomaches that were not completely filled after the soup, were topped off with peanut butter and/or jelly tortillas.

The PM-EVA was started not long after lunch. Nicky, Victoria and Aster suited up and could enjoy warmer weather than the first EVA. They didn't have any fogging problems at all. Victoria took some snow samples, Nicky took soil and water samples. Aster tried documenting the EVA with a small helmet camera.

With all EVA's done, the next big point on the planning was supper. Chili tortillas with rehydrated broccoli were the culinary highlight of the day! Since we are participating in a food study, we need to fill in daily surveys. So while Leandro's brownies are getting ready in the oven for desert, that will be our last homework for today. 


Chef's Report

Date: 18 December 2011

Written by Aster Stein and Nicky De Munster of Crew 109

1a) Today was

- a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) or activities were prepared to mark the celebration?

NO

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Premixed chili, dehydrated beef cubes, dehydrated broccoli, kool-aid, brownies

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

2 servings of chili, 1 serving of broccoli, 3 servings of kool-aid

4) Recipes for anything more complicated than following package directions exactly, or rehydrating. 

None so far

5) Lessons learned

Not yet

6) Comments/questions for Kim and Jean, the food study investigators

Our kitchen cupboards are filled with lots of salt, mustard, baking powder and baking soda

- any problems with supplied food?

no

- any problems with kitchen equipment, including cleaning supplies?

no

- any cooking questions you need help with?

The difference between cooking and non-cooking isn’t 100% clear

- any unusual events? (cleaned the oven, spilled all the vanilla, took inventory)

no

- any other thoughts or suggestions?

no